# Caffarelli’s work on elliptic free boundary problems

### v1

@@ 8. The $\varepsilon$-regularity theorem
The latter is identically 00 on $H_{0}$ and by a Schwartz reflection we can extend it harmonically to all of $\mathbb{R}^{n}$. The quadratic lower bound and Liouville’s theorem imply that $h$ is a second order polynomial, but in fact it has to be homoegenous and quadratic because $h_{0,0}(0)=0$ and $\nabla h_{0,0}(0)=0$. However a quadratic polynomial which vanishes on the hyperplane $H_{0}$ is necessarily a function of the variable $x_{n}$ only and, because of the harmonicity, this forces $h$ to vanish identically. So there is a unique blow-up of $v_{0}$ at the origin and it is the function
@@ 8.1. Lipschitz regularity for $\partial\{v_{0}>0\}$
In the first step we will derive the Lipschitz regularity of $\partial\{v_{0}>0\}$ in a neighborhood of 00 using a “PDE argument”. Even though we already know this from the convexity of $\{v_{0}=0\}$, the argument will in fact be functional to proving Theorem 8.3 because we will be able to export it to the “perturbed” setting of the theorem.
@@ 8.2. $C^{1,\alpha}$ regularity of $\partial\{v_{0}>0\}$
is Hölder continuous. Recall that indeed $\frac{\partial v_{0}}{\partial y_{2}}$ is positive. In fact the ratio is bounded away from zero (again by the Harnack inequality). So the inverse of the ratio in (8.3) is as well Hölder and, because the product of Hölder functions is Hölder, we finally conclude that $\frac{\nabla v_{0}}{|\nabla v_{0}|}$ is Hölder.
We now can invoke the same argument of the previous section: the vector field $\frac{\nabla v_{0}}{|\nabla v_{0}|}$ gives in fact the unit normal to the level sets of $v_{0}$ on $\{v_{0}>0\}$ and its Hölder continuity means the Hölder continuity of the tangent plane to such level sets. Since our estimate is however independent of which level set we are looking at, it can be transferred to $\partial\{v_{0}=0\}$ (as we transferred the Lipschitz bound).
@@
contradicting $B_{1/2}(x_{0})\subset\{v_{0,0}=0\}$. In fact the above argument, suitably modified, proves the following
@@
where $\delta$ is again a nonnegative number which will be chosen appropriately in few lines. Now $k(x_{0})=h(x_{0})-\delta v(x_{0})<0$. By the maximum principle $v$ must have a negative minimum on $\partial(B_{1/4}(x_{0})\cap\Omega)$. The latter minimum cannot however be on $\partial\Omega$, because there both $h$ and $v$ vanish and hence $k$ must be positive. Thus
@@ 9. Teaser
As already mentioned in the introduction, the “blow-up and partial regularity” program has been carried by Luis and some other coauthors in all the four problems mentioned in Section 3. We have examined in fairly many details the first of the problems, for the other three the starting points are the papers [4, 1, 2]. The theory developed by Luis and coauthors cover much more general situations than the model ones mentioned in Section 3 and many results extend to the parabolic setting. For an account of many general aspects we refer the interested reader to the book [6].

### v2

@@ 8. The $\varepsilon$-regularity theorem
The latter is identically 00 on $H_{0}$ and by a Schwartz reflection we can extend it harmonically to all of $\mathbb{R}^{n}$. The quadratic lower bound and Liouville’s theorem imply that $h$ is a second order polynomial, but in fact it has to be homogenous and quadratic because $h_{0,0}(0)=0$ and $\nabla h_{0,0}(0)=0$. However a quadratic polynomial which vanishes on the hyperplane $H_{0}$ is necessarily a function of the variable $x_{n}$ only and, because of the harmonicity, this forces $h$ to vanish identically. So there is a unique blow-up of $v_{0}$ at the origin and it is the function
@@ 8.1. Lipschitz regularity for $\partial\{v_{0}>0\}$
In the first step we will derive the Lipschitz regularity of $\partial\{v_{0}>0\}$ in a neighborhood of 00 using a “PDE argument”. Even though we already know this from the convexity of $\{v_{0}=0\}$, the argument will in fact be useful to prove Theorem 8.3 because we will be able to export it to the “perturbed” setting of the theorem.
@@ 8.2. $C^{1,\alpha}$ regularity of $\partial\{v_{0}>0\}$
is Hölder continuous. Recall that indeed $\frac{\partial v_{0}}{\partial y_{2}}$ is positive. In fact the ratio is bounded away from zero (again by the boundary Harnack inequality). So the inverse of the ratio in (8.3) is as well Hölder and, because the product of Hölder functions is Hölder, we finally conclude that $\frac{\nabla v_{0}}{|\nabla v_{0}|}$ is Hölder.
We now can invoke the same argument of the previous section: the vector field $\frac{\nabla v_{0}}{|\nabla v_{0}|}$ gives in fact the unit normal to the level sets of $v_{0}$ on $\{v_{0}>0\}$ and its Hölder continuity means the Hölder continuity of the tangent plane to such level sets. Since our estimate is independent of which level set we are looking at, it can be transferred to $\partial\{v_{0}=0\}$ (as we transferred the Lipschitz bound).
@@
contradicting $B_{1/2}(x_{0})\subset\{v_{0,0}=0\}$. In fact, the above argument, suitably modified, proves the following
@@
where $\delta$ is again a nonnegative number which will be chosen appropriately in a few lines. Now $k(x_{0})=h(x_{0})-\delta v(x_{0})<0$. By the maximum principle $v$ must have a negative minimum on $\partial(B_{1/4}(x_{0})\cap\Omega)$. The latter minimum cannot be on $\partial\Omega$, because there both $h$ and $v$ vanish and hence $k$ must be positive. Thus
@@ 9. Teaser
As already mentioned in the introduction, the “blow-up and partial regularity” program has been carried by Luis and some other co-authors in all the four problems mentioned in Section 3. We have examined in fairly many details the first of the problems, for the other three the starting points are the papers [4, 1, 2]. The theory developed by Luis and co-authors covers much more general situations than the model ones mentioned in Section 3 and many results extend to the parabolic setting. For an account of many general aspects we refer the interested reader to the book [6].